# Reassessing quantum-thermodynamic enhancements in continuous thermal machines

### v2

@@ D.2 Three-level coherent amplifier
$\begin{split}&a_{0}^{\prime}=2g^{2}\gamma_{c}\gamma_{h}\left(\bar{n}_{c}-\bar{n}_{h}\right)\\ &a_{0}^{\prime\prime}=-2g^{2}\gamma_{c}\gamma_{h}\left(\bar{n}_{c}\left(2\bar{n}_{h}+1\right)+\bar{n}_{h}\right)\\ &a_{1}=\frac{1}{2}\left(\gamma_{c}\left(\bar{n}_{c}\left(\gamma_{h}^{2}\bar{n}_{h}\left(3\bar{n}_{h}+1\right)+12g^{2}\right)+\gamma_{h}^{2}\bar{n}_{h}^{2}+8g^{2}\right)+\gamma_{c}^{2}\gamma_{h}\bar{n}_{c}\left(\bar{n}_{c}\left(3\bar{n}_{h}+1\right)+\bar{n}_{h}\right)+4g^{2}\gamma_{h}\left(3\bar{n}_{h}+2\right)\right)\\ &a_{2}=\frac{1}{2}\left(\gamma_{c}\gamma_{h}\left(\bar{n}_{c}\left(10\bar{n}_{h}+3\right)+3\bar{n}_{h}\right)+\gamma_{c}^{2}\bar{n}_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}^{2}\bar{n}_{h}\left(2\bar{n}_{h}+1\right)+8g^{2}\right).\end{split}$ (78)
and $a_{1}^{\prime}=0$.
In the case of the classical equivalent, we no longer consider the contribution of coherence to the dynamics, but introduce the classical stochastic transition rate between the interacting levels, leading to a matrix $W_{G}^{\mathrm{cl}}$ that reads:
$W_{G}^{\mathrm{cl}}=\left(\begin{array}[]{ccc}-(\gamma_{02}-\gamma_{10}^{\mathrm{cl}})&\gamma_{10}^{\mathrm{cl}}&\gamma_{20}~{}\mathrm{e}^{-i\chi_{\mathrm{h}}}\\ \gamma_{10}^{\mathrm{cl}}&-(\gamma_{12}+\gamma_{10}^{\mathrm{cl}})&\gamma_{21}~{}\mathrm{e}^{-i\chi_{\mathrm{c}}}\\ \gamma_{02}~{}\mathrm{e}^{i\chi_{\mathrm{h}}}&\gamma_{12}~{}\mathrm{e}^{i\chi_{\mathrm{c}}}&-(\gamma_{20}+\gamma_{21})\end{array}\right),\$
@@
$\begin{split}&a_{0}^{\mathrm{cl}\,\prime}=\frac{4g^{2}\gamma_{c}\gamma_{h}\left(\bar{n}_{h}-\bar{n}_{c}\right)}{\gamma_{c}\bar{n}_{c}+\gamma_{h}\bar{n}_{h}}\\ &a_{0}^{\mathrm{cl}\,\prime\prime}=\frac{4g^{2}\gamma_{c}\gamma_{h}\left(\bar{n}_{h}-\bar{n}_{c}\right)}{\gamma_{c}\bar{n}_{c}+\gamma_{h}\bar{n}_{h}}\\ &a_{1}^{\mathrm{cl}}=-\frac{8g^{2}\left(\gamma_{c}+\gamma_{h}\right)}{\gamma_{c}\bar{n}_{c}+\gamma_{h}\bar{n}_{h}}-\gamma_{c}\gamma_{h}\bar{n}_{c}-\gamma_{c}\gamma_{h}\left(3\bar{n}_{c}+1\right)\bar{n}_{h}-12g^{2}\\ &a_{2}^{\mathrm{cl}}=-\frac{8g^{2}}{\gamma_{c}\bar{n}_{c}+\gamma_{h}\bar{n}_{h}}-\gamma_{c}\left(2\bar{n}_{c}+1\right)-\gamma_{h}\left(2\bar{n}_{h}+1\right).\end{split}$ (79)
and again $a_{1}^{\mathrm{cl}\,\prime}=0$.
D.3 Three-qubit autonomous refrigerator
@@
Using the Hamiltonian and the Lindblad operators given in Sec. II.2, we can derive in an analogous way the $W_{G}$ matrices for the autonomous absorption refrigerator and its classical equivalent. Here again, we consider only relevant elements of the density matrix, which now consist of the populations of the $8$ energy levels (in the three-qubit composed ladder) and a pure imaginary coherence between the states |$\ket{101}$ and |$\ket{010}$. The $W_{G}$ matrix now have the form:
@@
$\begin{split}&W_{G}=\left(\begin{array}[]{cccccccccc}\Gamma_{1}&\gamma_{10}^{\mathrm{h}}\mathrm{e}^{-i\chi_{\mathrm{h}}}&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0&0&0&0&\gamma_{10}^{\mathrm{m}}\mathrm{e}^{-i\chi_{\mathrm{m}}}&0\\ \gamma_{01}^{\mathrm{h}}\mathrm{e}^{i\chi_{\mathrm{h}}}&\Gamma_{2}&0&\gamma_{10}^{\mathrm{m}}\mathrm{e}^{-i\chi_{\mathrm{m}}}&0&0&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0&0\\ \gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi_{\mathrm{c}}}&0&\Gamma_{3}&0&\gamma_{10}^{\mathrm{m}}\mathrm{e}^{-i\chi_{\mathrm{m}}}&0&\gamma_{10}^{\mathrm{h}}\mathrm{e}^{-i\chi_{\mathrm{h}}}&0&0\\ 0&\gamma_{01}^{\mathrm{m}}\mathrm{e}^{i\chi_{\mathrm{m}}}&0&\Gamma_{4}&0&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0&\gamma_{01}^{\mathrm{h}}\mathrm{e}^{i\chi_{\mathrm{h}}}&0\\ 0&0&\gamma_{01}^{\mathrm{m}}\mathrm{e}^{i\chi_{\mathrm{m}}}&0&\Gamma_{5}&\gamma_{10}^{\mathrm{h}}\mathrm{e}^{-i\chi_{\mathrm{h}}}&0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi_{\mathrm{c}}}&0\\ 0&0&0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi_{\mathrm{c}}}&\gamma_{01}^{\mathrm{h}}\mathrm{e}^{i\chi_{\mathrm{h}}}&\Gamma_{6}&\gamma_{01}^{\mathrm{m}}\mathrm{e}^{i\chi_{\mathrm{m}}}&0&0\\ 0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi_{\mathrm{c}}}&\gamma_{01}^{\mathrm{h}}\mathrm{e}^{i\chi_{\mathrm{h}}}&0&0&\gamma_{10}^{\mathrm{m}}\mathrm{e}^{-i\chi_{\mathrm{m}}}&\Gamma_{7}&0&g\\ \gamma_{01}^{\mathrm{m}}\mathrm{e}^{i\chi_{\mathrm{m}}}&0&0&\gamma_{10}^{\mathrm{h}}\mathrm{e}^{-i\chi_{\mathrm{h}}}&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0&0&\Gamma_{8}&-g\\ 0&0&0&0&0&0&-g&g&\Gamma_{9}\\ \end{array}\right)\end{split}$
@@
associated to vector $\vec{p}_{G}(t)=(\rho_{000},\rho_{001},\rho_{100},\rho_{011},\rho_{111},\rho_{101},\rho_{010},\mathrm{Im}[\rho_{010-101}])$. Some relevant inverse counting coefficients are in this case ((all coefficients can be found in [97])):
$\begin{split}&a_{0}^{\prime}=2g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}+\bar{n}_{m}+1\right)-\bar{n}_{c}\bar{n}_{m}\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\\ &\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\\ &a_{0}^{\prime\prime}=2g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}\left(2\bar{n}_{m}+1\right)+\bar{n}_{m}+1\right)+\bar{n}_{c}\bar{n}_{m}\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\\ &\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\\ &a_{1}^{\prime}=4g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}+\bar{n}_{m}+1\right)-\bar{n}_{c}\bar{n}_{m}\right)(4\gamma_{c}^{2}\left(2\bar{n}_{c}+1\right){}^{2}\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)+\gamma_{c}\left(2\bar{n}_{c}+1\right)(9\gamma_{h}\gamma_{m}\left(2\bar{n}_{h}+1\right)\\ &\left(2\bar{n}_{m}+1\right)+4\gamma_{h}^{2}\left(2\bar{n}_{h}+1\right){}^{2}+4\gamma_{m}^{2}\left(2\bar{n}_{m}+1\right){}^{2})+\gamma_{c}^{3}\left(2\bar{n}_{c}+1\right){}^{3}+\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)(3\gamma_{h}\gamma_{m}\left(2\bar{n}_{h}+1\right)\\ &\left(2\bar{n}_{m}+1\right)+\gamma_{h}^{2}\left(2\bar{n}_{h}+1\right){}^{2}+\gamma_{m}^{2}\left(2\bar{n}_{m}+1\right){}^{2}))\\ \end{split}$ (80)
@@
$\begin{split}&W_{G}^{\mathrm{cl}}=\left(\begin{array}[]{ccccccccc}\Gamma_{1}&\gamma_{10}^{\mathrm{h}}\mathrm{e}^{-i\chi_{\mathrm{h}}}&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0&0&0&0&\gamma_{10}^{\mathrm{m}}\mathrm{e}^{-i\chi_{\mathrm{m}}}\\ \gamma_{01}^{\mathrm{h}}\mathrm{e}^{i\chi_{\mathrm{h}}}&\Gamma_{2}&0&\gamma_{10}^{\mathrm{m}}\mathrm{e}^{-i\chi_{\mathrm{m}}}&0&0&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0\\ \gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi_{\mathrm{c}}}&0&\Gamma_{3}&0&\gamma_{10}^{\mathrm{m}}\mathrm{e}^{-i\chi_{\mathrm{m}}}&0&\gamma_{10}^{\mathrm{h}}\mathrm{e}^{-i\chi_{\mathrm{h}}}&0\\ 0&\gamma_{01}^{\mathrm{m}}\mathrm{e}^{i\chi_{\mathrm{m}}}&0&\Gamma_{4}&0&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0&\gamma_{01}^{\mathrm{h}}\mathrm{e}^{i\chi_{\mathrm{h}}}\\ 0&0&\gamma_{01}^{\mathrm{m}}\mathrm{e}^{i\chi_{\mathrm{m}}}&0&\Gamma_{5}&\gamma_{10}^{\mathrm{h}}\mathrm{e}^{-i\chi_{\mathrm{h}}}&0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi_{\mathrm{c}}}\\ 0&0&0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi_{\mathrm{c}}}&\gamma_{01}^{\mathrm{h}}\mathrm{e}^{i\chi_{\mathrm{h}}}&\Gamma_{6}&\gamma_{01}^{\mathrm{m}}\mathrm{e}^{i\chi_{\mathrm{m}}}&0\\ 0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi_{\mathrm{c}}}&\gamma_{01}^{\mathrm{h}}\mathrm{e}^{i\chi_{\mathrm{h}}}&0&0&\gamma_{10}^{\mathrm{m}}\mathrm{e}^{-i\chi_{\mathrm{m}}}&\Gamma_{7}-\gamma^{\mathrm{cl}}&\gamma^{\mathrm{cl}}\\ \gamma_{01}^{\mathrm{m}}\mathrm{e}^{i\chi_{\mathrm{m}}}&0&0&\gamma_{10}^{\mathrm{h}}\mathrm{e}^{-i\chi_{\mathrm{h}}}&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0&\gamma^{\mathrm{cl}}&\Gamma_{8}-\gamma^{\mathrm{cl}}\\ \end{array}\right)\end{split}$

### v4

@@ F.2 Three-level coherent amplifier
$\begin{split}&a_{0}^{\prime}=2g^{2}\gamma_{c}\gamma_{h}\left(\bar{n}_{c}-\bar{n}_{h}\right),~{}~{}~{}~{}a_{0}^{\prime\prime}=-2g^{2}\gamma_{c}\gamma_{h}\left(\bar{n}_{c}\left(2\bar{n}_{h}+1\right)+\bar{n}_{h}\right),\\ &a_{1}=\frac{1}{2}\left(\gamma_{c}\left(\bar{n}_{c}\left(\gamma_{h}^{2}\bar{n}_{h}\left(3\bar{n}_{h}+1\right)+12g^{2}\right)+\gamma_{h}^{2}\bar{n}_{h}^{2}+8g^{2}\right)+\gamma_{c}^{2}\gamma_{h}\bar{n}_{c}\left(\bar{n}_{c}\left(3\bar{n}_{h}+1\right)+\bar{n}_{h}\right)+4g^{2}\gamma_{h}\left(3\bar{n}_{h}+2\right)\right),\\ &a_{2}=\frac{1}{2}\left(\gamma_{c}\gamma_{h}\left(\bar{n}_{c}\left(10\bar{n}_{h}+3\right)+3\bar{n}_{h}\right)+\gamma_{c}^{2}\bar{n}_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}^{2}\bar{n}_{h}\left(2\bar{n}_{h}+1\right)+8g^{2}\right).\end{split}$ (79)
and $a_{1}^{\prime}=0$. In the case of the classical equivalent, we no longer consider the contribution of coherence to the dynamics, but introduce the classical stochastic transition rate between the interacting levels, leading to a matrix $W_{G}^{\mathrm{cl}}$ that reads:
$W_{G}^{\mathrm{cl}}=\left(\begin{array}[]{ccc}-(\gamma_{02}-\gamma_{10}^{\mathrm{cl}})&\gamma_{10}^{\mathrm{cl}}&\gamma_{20}\\ \gamma_{10}^{\mathrm{cl}}&-(\gamma_{12}+\gamma_{10}^{\mathrm{cl}})&\gamma_{21}~{}\mathrm{e}^{-i\chi}\\ \gamma_{02}&\gamma_{12}~{}\mathrm{e}^{i\chi}&-(\gamma_{20}+\gamma_{21})\end{array}\right),\$
@@
$a_{0}^{\mathrm{cl}\,\prime}=\frac{4g^{2}\gamma_{c}\gamma_{h}\left(\bar{n}_{h}-\bar{n}_{c}\right)}{\gamma_{c}\bar{n}_{c}+\gamma_{h}\bar{n}_{h}},~{}~{}~{}~{}a_{0}^{\mathrm{cl}\,\prime\prime}=\frac{4g^{2}\gamma_{c}\gamma_{h}\left(\bar{n}_{h}-\bar{n}_{c}\right)}{\gamma_{c}\bar{n}_{c}+\gamma_{h}\bar{n}_{h}},~{}~{}~{}~{}a_{1}^{\mathrm{cl}\,\prime}=0,$ (80)
$a_{1}^{\mathrm{cl}}=-\frac{8g^{2}\left(\gamma_{c}+\gamma_{h}\right)}{\gamma_{c}\bar{n}_{c}+\gamma_{h}\bar{n}_{h}}-\gamma_{c}\gamma_{h}\bar{n}_{c}-\gamma_{c}\gamma_{h}\left(3\bar{n}_{c}+1\right)\bar{n}_{h}-12g^{2},~{}~{}~{}~{}a_{2}^{\mathrm{cl}}=-\frac{8g^{2}}{\gamma_{c}\bar{n}_{c}+\gamma_{h}\bar{n}_{h}}-\gamma_{c}\left(2\bar{n}_{c}+1\right)-\gamma_{h}\left(2\bar{n}_{h}+1\right).$
F.3 Three-qubit autonomous refrigerator
@@
Using the Hamiltonian and the Lindblad operators given in Eq. (29) and (30), we can derive in an analogous way the $W_{G}$ matrices for the autonomous absorption refrigerator and its classical equivalent. Here again, we consider only relevant elements of the density matrix, which now consist of the populations of the $8$ energy levels (in the three-qubit composed ladder) and a pure imaginary coherence between the states |$\ket{101}$ and |$\ket{010}$. The $W_{G}$ matrix now have the form:
@@
$\begin{split}&W_{G}=\left(\begin{array}[]{cccccccccc}\Gamma_{1}&\gamma_{10}^{\mathrm{h}}&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi}&0&0&0&0&\gamma_{10}^{\mathrm{m}}&0\\ \gamma_{01}^{\mathrm{h}}&\Gamma_{2}&0&\gamma_{10}^{\mathrm{m}}&0&0&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi}&0&0\\ \gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi}&0&\Gamma_{3}&0&\gamma_{10}^{\mathrm{m}}&0&\gamma_{10}^{\mathrm{h}}&0&0\\ 0&\gamma_{01}^{\mathrm{m}}&0&\Gamma_{4}&0&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi}&0&\gamma_{01}^{\mathrm{h}}&0\\ 0&0&\gamma_{01}^{\mathrm{m}}&0&\Gamma_{5}&\gamma_{10}^{\mathrm{h}}&0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi}&0\\ 0&0&0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi}&\gamma_{01}^{\mathrm{h}}&\Gamma_{6}&\gamma_{01}^{\mathrm{m}}&0&0\\ 0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi}&\gamma_{01}^{\mathrm{h}}&0&0&\gamma_{10}^{\mathrm{m}}&\Gamma_{7}&0&g\\ \gamma_{01}^{\mathrm{m}}&0&0&\gamma_{10}^{\mathrm{h}}&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi}&0&0&\Gamma_{8}&-g\\ 0&0&0&0&0&0&-g&g&\Gamma_{9}\\ \end{array}\right),\end{split}$
associated to vector $\vec{p}_{G}(t)=(\rho_{000},\rho_{001},\rho_{100},\rho_{011},\rho_{111},\rho_{101},\rho_{010},\mathrm{Im}[\rho_{010-101}])$. In the above equation, we defined the aggregated rates $\Gamma_{1}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$, $\Gamma_{2}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{3}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$, $\Gamma_{4}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{5}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$, $\Gamma_{6}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{7}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{8}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$ and $\Gamma_{9}=-\frac{1}{2}(\gamma_{d}+\gamma_{2d}+\gamma_{3d}+\gamma_{u}+\gamma_{2u}+\gamma_{3u})$.
@@
Some relevant inverse counting coefficients are in this case (all coefficients can be found in [92]):
$a_{0}^{\prime}=$ $2g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}+\bar{n}_{m}+1\right)-\bar{n}_{c}\bar{n}_{m}\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)$
$\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right),$ (81)
$a_{0}^{\prime\prime}=$ $2g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}\left(2\bar{n}_{m}+1\right)+\bar{n}_{m}+1\right)+\bar{n}_{c}\bar{n}_{m}\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)$
$\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right),$
$a_{1}^{\prime}=$ $4g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}+\bar{n}_{m}+1\right)-\bar{n}_{c}\bar{n}_{m}\right)(4\gamma_{c}^{2}\left(2\bar{n}_{c}+1\right){}^{2}\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)+\gamma_{c}\left(2\bar{n}_{c}+1\right)(9\gamma_{h}\gamma_{m}\left(2\bar{n}_{h}+1\right)$
$\left(2\bar{n}_{m}+1\right)+4\gamma_{h}^{2}\left(2\bar{n}_{h}+1\right){}^{2}+4\gamma_{m}^{2}\left(2\bar{n}_{m}+1\right){}^{2})+\gamma_{c}^{3}\left(2\bar{n}_{c}+1\right){}^{3}+\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)(3\gamma_{h}\gamma_{m}\left(2\bar{n}_{h}+1\right)$
$\left(2\bar{n}_{m}+1\right)+\gamma_{h}^{2}\left(2\bar{n}_{h}+1\right){}^{2}+\gamma_{m}^{2}\left(2\bar{n}_{m}+1\right){}^{2})).$
@@
$\begin{split}&W_{G}^{\mathrm{cl}}=\left(\begin{array}[]{ccccccccc}\Gamma_{1}&\gamma_{10}^{\mathrm{h}}&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi}&0&0&0&0&\gamma_{10}^{\mathrm{m}}\\ \gamma_{01}^{\mathrm{h}}&\Gamma_{2}&0&\gamma_{10}^{\mathrm{m}}&0&0&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi}&0\\ \gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi}&0&\Gamma_{3}&0&\gamma_{10}^{\mathrm{m}}&0&\gamma_{10}^{\mathrm{h}}&0\\ 0&\gamma_{01}^{\mathrm{m}}&0&\Gamma_{4}&0&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi}&0&\gamma_{01}^{\mathrm{h}}\\ 0&0&\gamma_{01}^{\mathrm{m}}&0&\Gamma_{5}&\gamma_{10}^{\mathrm{h}}&0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi}\\ 0&0&0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi}&\gamma_{01}^{\mathrm{h}}&\Gamma_{6}&\gamma_{01}^{\mathrm{m}}&0\\ 0&\gamma_{01}^{\mathrm{c}}\mathrm{e}^{i\chi}&\gamma_{01}^{\mathrm{h}}&0&0&\gamma_{10}^{\mathrm{m}}&\Gamma_{7}-\gamma^{\mathrm{cl}}&\gamma^{\mathrm{cl}}\\ \gamma_{01}^{\mathrm{m}}&0&0&\gamma_{10}^{\mathrm{h}}&\gamma_{10}^{\mathrm{c}}\mathrm{e}^{-i\chi}&0&\gamma^{\mathrm{cl}}&\Gamma_{8}-\gamma^{\mathrm{cl}}\\ \end{array}\right),\end{split}$